Determine the pH, the pOH, the [OH-] and [H+] of a 0.025M barium hydroxide solution.
Answer #1

2+ - BaloH) I Batoh pOH = -log [oH] = -log (0.025) = 1.6020 PH+ POH = 14 pH = 14 - POH = 14-1.6020 = 12.398 .pH = -log [ht ]